# Improving accuracy of tree-tensor network approach by optimization of network structure

### v1

@@ Appendix C Exact solution of Richardson model
where $S^{\pm}_{i}=S^{x}_{i}\pm iS^{y}_{i}$, and $b_{i}$ is the hardcore-boson operator at $i$ th site obeying $\left[b_{i},b_{i^{\prime}}^{\dagger}\right]=\left(1-2b_{i}^{\dagger}b_{i}\right)\delta_{ii^{\prime}}$ and $\left(b_{i}^{\dagger}\right)^{2}=0$. Substituting the expressions of Eq. (18) into the spin Hamiltonian $\mathcal{H}_{\rm R}$, one can map the spin model into the interacting hardcore-boson model ~$\tilde{\mathcal{H}}_{\rm R}$ as
$\mathcal{H}_{\rm R}$ $=$ ~$\tilde{\mathcal{H}}_{\rm R}-\frac{1}{2}\sum_{i=1}^{N}h_{i}+gN_{\rm b},$ (19)
~$\tilde{\mathcal{H}}_{\rm R}$ $=$ $\sum_{i=1}^{N}\sum_{i^{\prime}=1}^{N}\left(h_{i}\delta_{ii^{\prime}}-g\right)b^{\dagger}_{i}b_{i^{\prime}}.$ (20)
@@
The eigenstates of the Hamiltonian ~$\tilde{\mathcal{H}}_{\rm R}$ [Eq. (20)] can be written down explicitly as
@@
where $|0\rangle$ is the vacuum of the bosons $\{b_{i}\}$. The state $|\tilde{\Psi}_{\rm R}\rangle$ in Eq. (23) is an exact eigenstate of the Hamiltonian ~$\tilde{\mathcal{H}}_{\rm R}$ if the parameters $\{E_{n}\}$ satisfy the coupled equations,
@@
Note that in these equations, $y_{p}$ appears only in the form of $y_{p}^{2}$. We thus treat $\{x_{p}\}$ and $\{y_{p}^{2}\}$ as real parameters and solve the coupled equations (28) and (29) numerically using the Newton-Raphson method[46].
@@
Here, note that the field $h_{i}$ is arranged in the ascending order. Furthermore, for $g=-\delta g$ with a small $\delta g>0$, the approximate ground-state solution up to the first order of $\delta g$ is obtained as
@@
and input them into the Newton-Raphson method. In such a manner, we obtained the ground-state solutions as a function of $g$ with decreasing $g$ [increasing $J_{\rm R}$ in Eq. (21)] adiabatically. The obtained ground-state energies for the model parameters discussed in Sec. III.3 are presented in Table 2.

### v2

@@ Appendix D Exact solution of Richardson model
where $S^{\pm}_{i}=S^{x}_{i}\pm iS^{y}_{i}$, and $b_{i}$ is the hardcore-boson operator at the $i$ th site obeying $\left[b_{i},b_{i^{\prime}}^{\dagger}\right]=\left(1-2b_{i}^{\dagger}b_{i}\right)\delta_{ii^{\prime}}$ and $\left(b_{i}^{\dagger}\right)^{2}=0$. Substituting the expressions of Eq. (18) into the spin Hamiltonian $\mathcal{H}_{\rm R}$, one can map the spin model into the interacting hardcore-boson model $\tilde{\mathcal{H}}_{\rm R}$ as
$\mathcal{H}_{\rm R}$ $=$ $\tilde{\mathcal{H}}_{\rm R}-\frac{1}{2}\sum_{i=1}^{N}h_{i}+gN_{\rm b},$ (19)
$\tilde{\mathcal{H}}_{\rm R}$ $=$ $\sum_{i=1}^{N}\sum_{i^{\prime}=1}^{N}\left(h_{i}\delta_{ii^{\prime}}-g\right)b^{\dagger}_{i}b_{i^{\prime}}.$ (20)
@@
The eigenstates of the Hamiltonian $\tilde{\mathcal{H}}_{\rm R}$ [Eq. (20)] can be written down explicitly as
@@
where $|0\rangle$ is the vacuum of the bosons $\{b_{i}\}$. The state $|\tilde{\Psi}_{\rm R}\rangle$ in Eq. (23) is an exact eigenstate of the Hamiltonian $\tilde{\mathcal{H}}_{\rm R}$ if the parameters $\{E_{n}\}$ satisfy the coupled equations,
@@
Note that in these equations, $y_{p}$ appears only in the form of $y_{p}^{2}$. Therefore, we can treat $\{x_{p}\}$ and $\{y_{p}^{2}\}$ as real parameters and solve the coupled equations (28) and (29) numerically using the Newton-Raphson method[49].
@@
Here, note that the field $h_{i}$ is arranged in ascending order. Furthermore, for $g=-\delta g$ with a small $\delta g>0$, the approximate ground-state solution up to the first order of $\delta g$ is obtained as
@@
and input them into the Newton-Raphson method. In such a manner, we obtained the ground-state solutions as a function of $g$ by decreasing $g$ [increasing $J_{\rm R}$ in Eq. (21)] adiabatically. The obtained ground-state energies for the model parameters discussed in Sec. III.3 are presented in Table 2.